METCRAX: Logbook Entries

METCRAX: Site East_Lower Messages, 13 Entries..

Return to Logbook Contents Page
Entry Date Title Site Author #Graphics
164 Fri 03-Nov-2006el teardownEast_LowerOncley
138 Thu 26-Oct-2006logger change at ELEast_LowerSemmer
136 Wed 25-Oct-2006el logger died againEast_LowerOncley
133 Wed 25-Oct-2006TRH orderEast_LowerOncley
131 Wed 25-Oct-2006krypton kleanEast_LowerOncley
121 Tue 24-Oct-2006Battery change at logger ELEast_LowerSemmer
120 Tue 24-Oct-2006.5m sonic down at ELEast_LowerSemmer
70 Fri 13-Oct-2006el sensor serial numbersEast_LowerMartin
60 Thu 12-Oct-2006Tried to fix the RH at 3m elEast_LowerHorst
59 Thu 12-Oct-2006Fixed Tcase.out at elEast_LowerHorst
41 Sat 07-Oct-2006Tightened terminals in 50y data logger at euEast_LowerHorst
40 Sat 07-Oct-2006Tightened terminals in 50y data logger at elEast_LowerHorst
19 Fri 06-Oct-2006radiometers cleanedEast_LowerHoch

164: LOG, Site East_Lower, Fri 03-Nov-2006 21:11:38 MST, el teardown
Serial numbers from today's teardown of el:	16714	26416
PSP.out	17627
PYG.out	29260
Tsurf	3577-4
Rnet-parallel	Q97035
Rnet-horiz	20076
Gsoil	H013390	(pretty shallow -- only 3cm?)
Tsoil	STP97001
TP01	200235
logger	east lower	9864

[remaining numbers recorded by Kurt]

0.5m	CSAT	
1.5m	CSAT	
5m	CSAT	
8.5m	CSAT	

138: Radiometers, Site East_Lower, Thu 26-Oct-2006 14:22:47 MST, logger change at EL
The logger at EL now has a direct link to the ADAM. So far no drop out.
This occurred at ~12:45 local today.
136: Radiometers, Site East_Lower, Wed 25-Oct-2006 16:27:31 MST, el logger died again
died at about 1245 and still not back up.  The hard-wired logger does respond
to +++ commands, so the problem probably is on the logger side.  (Though the
logger battery was 13.2V and rising when it died.)

We're a bit puzzled.

133: TRH, Site East_Lower, Wed 25-Oct-2006 14:10:21 MST, TRH order
The TRH booms apparently were installed in the wrong order (despite being
clearly labeled).  We now know that the order of the sensors in the data
is: 1.5m, 5m, 8m, and 3m.  I will change the config now to reflect this

Steve played around with TRH.5m probe, but this did not recover the humidity 
131: KH2O, Site East_Lower, Wed 25-Oct-2006 14:07:08 MST, krypton klean
krypton kleaned ~1217.  MSB of counts went from 0e -> 1f.

121: Radiometers, Site East_Lower, Tue 24-Oct-2006 14:38:02 MST, Battery change at logger EL
Steve and I decided to change the logger battery at EL to see if it is causing
the logger outages during the middle of the day. This took place during a
crater visit between 11:30 and 13:00 local time.

120: Sonic, Site East_Lower, Tue 24-Oct-2006 10:21:10 MST, .5m sonic down at EL
The .5m sonic went down last night around 16:00 local time. Was able to get
it going again by logging in to LE, sr 5, and typing a "D".

70: Station_Visit, Site East_Lower, Fri 13-Oct-2006 17:31:33 MST, el sensor serial numbers
Q7		Q97035
Q7horiz		Q90076
pygin		EP26416
pygout		EP29260
pyrin		EP16714
pyrout		EP17877
Tsoil		STP97001
HFT		H013390
TP01		200235

60: TRH, Site East_Lower, Thu 12-Oct-2006 17:23:03 MST, Tried to fix the RH at 3m el
Swapped the wires for the 1.5m and 3m RH within the Campbell data logger and 
the bad data followed the 50Y.
Swapped the RH transducers between 1.5m and 3m and the problem stayed at 3m.
The conclusion is that the problem is in the wiring between the 50Y transducer 
and the Campbell data logger.
59: Radiometers, Site East_Lower, Thu 12-Oct-2006 17:21:37 MST, Fixed Tcase.out at el
Found a loose wire in the el Campbell data logger which was coming from Pyg.out.
Screwed it down and fixed the Tcase.out
41: TRH, Site East_Lower, Sat 07-Oct-2006 17:57:47 MST, Tightened terminals in 50y data logger at eu
Because 1.5 m RH is ratty at east upper, Charlie tightened the terminals in the 
50Y data logger.

40: TRH, Site East_Lower, Sat 07-Oct-2006 17:55:47 MST, Tightened terminals in 50y data logger at el
Because 3 m RH is ratty at east lower, Charlie tightened the terminals in the 
50Y data logger.
19: Radiometers, Site East_Lower, Fri 06-Oct-2006 10:46:44 MST, radiometers cleaned
03.10 15:15 cleaned 4 components, not net radiometers